# Anomalous dimensions from conformal field theory: generalized $\phi^{2n+1}$ theories

### v1

@@ 4 Generalized $\phi^{2n+1}$ theory ($n>1$)
This theory describes the generalized multicritical Yang-Lee edge singularity vonGehlen:1994rp ; Lencses:2022ira ; Lencses:2023evr ; Lencses:2024wib . For $k=1$, it was argued that the Lagrangian in (80) can be obtained from the Landau-Ginzburg Lagrangian of the multicritical Ising model. For the latter Lagrangian, a shift of the order parameter field by a purely imaginary constant generates $\mathbb{Z}_{2}$-odd terms. Then, one can tune the coupling constants to reach the multicritical point given by (80). [^24] Not all multicritical Ising models lead to the Lagrangian in (80). See Lencses:2024wib for more details. This generalizes Fisher’s argument in Fisher:1978pf .
@@ 4.1 Multiplet recombination
where $\gamma_{\phi}$ is given in (87). The special case (91) is associated with $\ell=m=0$. One can check that, to order $\alpha^{2}$, the anomalous dimension of the stress tensor $(m=k-1,\;\ell=2)$ vanishes, even though we do not know the explicit expression for $\alpha$. In other words, we cannot determine $\alpha$ by requiring that the stress tensor does not acquire anomalous dimension. [^25] We believe that the spinning recombination cannot fix $\alpha$ either. For example, at $k=1$, we find that $\lim_{\epsilon\rightarrow 0}(\alpha^{-2}\langle\Box\phi(x_{1})\phi(x_{2})\partial\mathcal{J}(x_{3}))\rangle$ is proportional to $\lim_{\epsilon\rightarrow 0}(\alpha^{-2}\gamma_{\mathcal{J}})$, so the matching condition should only yield a relation between $\gamma_{\mathcal{J}}$ and $\alpha$, and this relation ought to be the same as (95).
Since the anomalous dimensions are not fixed by our CFT analysis, we resort to inputs from the diagrammatic method. At $k=1$, Codello et al. computed the anomalous dimensions of some scalar operators in the $n>1$ theory Codello:2017epp . The result for $\gamma_{\phi}$ is expressed in terms of the critical coupling constant in their equation (5.7), which is consistent with our formula (87). Furthermore, they also obtained the critical coupling constant, so they fully determined the anomalous dimension $\gamma_{\phi}$. We fix $\alpha$ using their results and find [^26] In Codello:2017epp , the general expression for the critical coupling constant $g$ in equation (5.6) may contain a typo, as it does not seem to reproduce their $n=2$ result in equation (3.2). Here we use their beta functionals to solve for $g$, and the result is consistent with their equation (3.2).
@@
Here $\mathbb{Z}_{+}$ denotes the set of positive integers. If we use this input from the diagrammatic method, then the anomalous dimensions of $\phi$ and $\mathcal{J}$ are completely fixed for $k=1$. The $n=1$ case of (96) matches the Yang-Lee result (23). For $n=2,3,4$, our results for $\gamma_{\phi}$ are in agreement with those of Gracey:2017okb . [^27] However, our results for $\gamma_{\phi^{2}}$ do not agree with Gracey:2017okb . Some disagreement was pointed out by Codello:2017epp as well, and it was shown that the results of Codello:2017epp are supported by the CFT analysis Codello:2017qek .
@@ 4.2 Analytic bootstrap
In this subsection, we verify the multiplet-recombination results using the analytic bootstrap. For arbitrary $\alpha$, the multiplet-recombination results are consistent with crossing symmetry, according to the analytic bootstrap in this section. As in subsection 2.2, we consider the four-point function
@@
Figure 13: Feynman diagram in the Lagrangian description (80). Due to the $g\phi^{2n+1}$ vertices, the operator $\phi^{2n-1}$ contributes at order $g^{2}\sim\alpha^{2}$.
@@
The calculation of the dDisc associated with (4.2) amounts to multiplying by the factor $2\sin^{2}(\pi k\,\frac{2n-3}{2n-1})$. Substituting this dDisc into the inversion integral (102) and dividing by the free OPE coefficient (100), we obtain the anomalous dimensions of the bilinear operators on the $m=0$ trajectory. The results agree with those from the multiplet recombination, confirming that the multiplet-recombination results are crossing symmetric. In appendix B, the analytic bootstrap has also been carried out at subleading twist, where the multiplet-recombination results are again consistent with crossing symmetry.
@@ 5 Discussion
In this paper we have computed the anomalous dimensions of $\phi$ and bilinear operators $\mathcal{J}$ in the generalized $\phi^{2n+1}$ theory, using the conformal multiplet recombination. Our main results are summarized in table 1. We have verified that these results are consistent with crossing symmetry using the analytic bootstrap.
@@
As the generalized free theory is non-unitary, we expect that the $\phi^{2n+1}$ deformations are non-unitary as well. The nonperturbative bootstrap study of their integer $d$ versions requires bootstrap methods that do not rely on positivity constraints Gliozzi:2013ysa ; Gliozzi:2014jsa ; El-Showk:2016mxr ; Esterlis:2016psv ; Li:2017agi ; Li:2017ukc ; Li:2021uki ; Kantor:2021kbx ; Kantor:2021jpz ; Afkhami-Jeddi:2021iuw ; Laio:2022ayq ; Li:2023tic . In Guo:2023qtt and the present work, we have investigated the $\phi^{2n}$ and $\phi^{2n+1}$ deformations of the generalized free theory, i.e., $\phi^{m}$ interactions with integer power $m\geq 3$. A natural extension is to consider non-integer $m$. Some simple examples are the fractional power $m=\frac{m_{1}}{2},\frac{m_{1}}{3},\frac{m_{1}}{4},\dots$, where $m_{1}$ is a large enough integer. The $\phi^{m}$ interaction with irrational power or even complex power are also interesting from the perspective of analytic continuation, as they can connect the standard integer-power cases in the theory space, but we expect that they are non-unitary even for the canonical case of $k=1$. The $\mathcal{PT}$ symmetric theories are particularly relevant because they can satisfy the physical requirement of a bounded-from-below spectrum in a continuous range of power Bender:1998ke ; Bender:2007nj ; Bender:2023cem . Recently, the low dimensional $\mathcal{PT}$ symmetric models have been revisited by bootstrap methods in Li:2022prn ; Khan:2022uyz ; Li:2023nip ; John:2023him ; Li:2023ewe ; Li:2024rod , using the equations of motion in the Lagrangian or Hamiltonian formalism. It would be interesting to extend the low dimensional bootstrap methods to the higher dimensional cases. Along the lines of analytic continuation, one may also bootstrap the non-perturbative conformal field theory by the homotopy continuation method Li:2024ggr , where the $\epsilon$ expansion can provide regular boundary conditions near but away from the singular Gaussian limit.
@@ Appendix C Representation theory of $S_{N+1}$
$\dim V_{(N,1)}$ $=\frac{(N+1)!}{N\times(N-1)\times(N-3)\times\ldots\times 1\times 2\times 1}=\frac{(N+1)(N-2)}{2}\,,$ (130)
@@
Comparing the dimensions with those of (128), (129), (130) and (131), we conclude that (132), (133), (134) and (135) correspond to the Young diagrams $(N+1)$, $(N,1)$, $(N-1,2)$ and $(N-1,1^{2})$. [^30] The dimension of (134) follows from the symmetric tracelessness and the constraints $\sum_{a,b=1}^{N}T_{iab}\sum_{c,d=1}^{N}t^{(3)}_{abcd}\phi_{c}\phi_{d}=0\,,$ (137) where $i=1,2,\ldots,N$. From these examples, we see that the tensor product decomposition resembles that in the O($N$) model, except that here the symmetric traceless part further decomposes into two irreducible representations $V_{(N,1)}$ and $V_{(N-1,2)}$. Since $V_{(N,1)}$ is the standard representation, we use the term “symmetric traceless representation” to refer to $V_{(N-1,2)}$ below. Moreover, we refer to $V_{(N+1)}$ and $V_{(N-1,1^{2})}$ as the “trivial representation” and the “antisymmetric representation”. [^31] The antisymmetric representation $V_{(N-1,1^{2})}$ is the exterior power $\Lambda^{2}V_{(N,1)}$ of the standard representation $V_{(N,1)}$.
@@ Appendix D Tensor structures
The tensor product decomposition of $(N,1)^{4}$ contains the trivial representation four times, so there are four linearly independent rank-4 invariant tensors wallace1978spin . Using $e_{a}^{\alpha}$ defined by (126), we identify four rank-4 $S_{N+1}$-invariant tensors: $\delta_{ab}\delta_{cd}$, $\delta_{ac}\delta_{bd}$, $\delta_{ad}\delta_{bc}$ and $\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}$. We then use these tensors to define a basis for the four-point function $\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\phi_{c}(x_{3})\phi_{d}(x_{4})\rangle$:

### v2

@@ 4 Generalized $\phi^{2n+1}$ theory ($n>1$)
This theory describes the generalized multicritical Yang-Lee edge singularity vonGehlen:1994rp ; Lencses:2022ira ; Lencses:2023evr ; Lencses:2024wib . For $k=1$, it was argued that the Lagrangian in (80) can be obtained from the Landau-Ginzburg Lagrangian of the multicritical Ising model. For the latter Lagrangian, a shift of the order parameter field by a purely imaginary constant generates $\mathbb{Z}_{2}$-odd terms. Then, one can tune the coupling constants to reach the multicritical point given by (80).262626 This procedure can also lead to a multicritical point with an interaction term consisting of an even number of $\phi$’s. See Lencses:2024wib for more details. This generalizes Fisher’s argument in Fisher:1978pf .
@@ 4.1 Multiplet recombination
where $\gamma_{\phi}$ is given in (87). For $k=1$, this result can also be derived from eq.(4.19) in Gliozzi:2017gzh . The special case (91) is associated with $\ell=m=0$. One can check that, to order $\alpha^{2}$, the anomalous dimension of the stress tensor $(m=k-1,\;\ell=2)$ vanishes, even though we do not know the explicit expression for $\alpha$. In other words, we cannot determine $\alpha$ by requiring that the stress tensor does not acquire anomalous dimension. [^27] We believe that the spinning recombination cannot fix $\alpha$ either. For example, at $k=1$, we find that $\lim_{\epsilon\rightarrow 0}(\alpha^{-2}\langle\Box\phi(x_{1})\phi(x_{2})\partial\mathcal{J}(x_{3}))\rangle$ is proportional to $\lim_{\epsilon\rightarrow 0}(\alpha^{-2}\gamma_{\mathcal{J}})$, so the matching condition should only yield a relation between $\gamma_{\mathcal{J}}$ and $\alpha$, and this relation ought to be the same as (95).
Since the anomalous dimensions are not fixed by our CFT analysis, we resort to inputs from the diagrammatic method. At $k=1$, Codello et al. computed the anomalous dimensions of some scalar operators in the $n>1$ theory Codello:2017epp . The result for $\gamma_{\phi}$ is expressed in terms of the critical coupling constant in their equation (5.7), which is consistent with our formula (87). Furthermore, they also obtained the critical coupling constant, so they fully determined the anomalous dimension $\gamma_{\phi}$. We use their results to fix $\alpha$  [^28] In Codello:2017epp , the general expression for the critical coupling constant $g$ in equation (5.6) may contain a typo, as it does not seem to reproduce their $n=2$ result in equation (3.2). Here we use their beta functionals to solve for $g$, and the result is consistent with their equation (3.2).
@@
Here $\mathbb{Z}_{+}$ denotes the set of positive integers. If we use this input from the diagrammatic method, then the anomalous dimensions of $\phi$ and $\mathcal{J}$ are completely fixed for $k=1$. For $n=1$, there is only one term in (4.1), which is associated with $r=s=t=1$. We verify that the $n=1$ case of (96) matches the Yang-Lee result (23). For $n=2,3,4$, our results for $\gamma_{\phi}$ are in agreement with those of Gracey:2017okb . [^29] However, our results for $\gamma_{\phi^{2}}$ do not agree with Gracey:2017okb . Some disagreement was pointed out by Codello:2017epp as well, and it was shown that the results of Codello:2017epp are supported by the CFT analysis Codello:2017qek .
@@ 4.2 Analytic bootstrap
In this subsection, we verify the multiplet-recombination results using the analytic bootstrap. The multiplet-recombination results are consistent with the leading crossing constraints for arbitrary $\alpha$. Therefore, as in the multiplet recombination approach, we are not able determine $\alpha$ by the leading CFT analysis. As in subsection 2.2, we consider the four-point function
@@
Figure 13: Feynman diagram in the Lagrangian description (80). Due to the $g\phi^{2n+1}$ vertices, the operator $\phi^{2n-1}$ contributes at order $g^{2}\sim\alpha^{2}$. This is similar to the case of the $\phi^{2n}$ theory discussed in Henriksson:2020jwk .
@@
The calculation of the dDisc associated with (4.2) amounts to multiplying by the factor $2\sin^{2}(\pi k\,\frac{2n-3}{2n-1})$. Substituting this dDisc into the inversion integral (102) and dividing by the free OPE coefficient (100), we obtain the anomalous dimensions of the bilinear operators on the $m=0$ trajectory. The results agree with those from the multiplet recombination, confirming that the multiplet-recombination results are compatible with OPE associativity. In appendix B, the analytic bootstrap has also been carried out at subleading twist, where the multiplet-recombination results are again consistent with crossing symmetry.
@@ 5 Discussion
In this paper we have computed the anomalous dimensions of $\phi$ and bilinear operators $\mathcal{J}$ in the generalized $\phi^{2n+1}$ theories, using the conformal multiplet recombination. For $n=1$, we have also considered the $\phi^{3}$ theory with global symmetry $S_{N+1}$. Our main results are summarized in table 1. We have verified that they are consistent with crossing symmetry using the analytic bootstrap. To the best of our knowledge, new results in the canonical case $k=1$ include the anomalous dimensions of spinning bilinear operators in the Potts model. In the generalized case with a higher-derivative kinetic term, i.e. $k>1$, all the results for the anomalous dimensions and OPE coefficients should be new.
@@
As the generalized free theory is non-unitary, we expect that the $\phi^{2n+1}$ deformations are non-unitary as well. The nonperturbative bootstrap study of their integer $d$ versions requires bootstrap methods that do not rely on positivity constraints Gliozzi:2013ysa ; Gliozzi:2014jsa ; El-Showk:2016mxr ; Esterlis:2016psv ; Li:2017agi ; Li:2017ukc ; Li:2021uki ; Kantor:2021kbx ; Kantor:2021jpz ; Afkhami-Jeddi:2021iuw ; Laio:2022ayq ; Li:2023tic . In Guo:2023qtt and the present work, we have investigated the $\phi^{2n}$ and $\phi^{2n+1}$ deformations of the generalized free theory, i.e., $\phi^{m}$ interactions with integer power $m\geq 3$.  [^30] Can the generalized $\phi^{m}$ theories be realized by lattice models? A natural extension is to consider non-integer $m$. Some simple examples are the fractional power $m=\frac{m_{1}}{2},\frac{m_{1}}{3},\frac{m_{1}}{4},\dots$, where $m_{1}$ is a large enough integer. The $\phi^{m}$ deformations with irrational power or even complex power are also interesting from the perspective of analytic continuation, as they can connect the standard integer-power cases in the theory space, but we expect that they are non-unitary even for the canonical case of $k=1$. The $\mathcal{PT}$ symmetric theories are particularly relevant because they can satisfy the physical requirement of a bounded-from-below spectrum in a continuous range of power Bender:1998ke ; Bender:2007nj ; Bender:2023cem .  [^31] It is not clear to us how to capture the $\mathcal{PT}$ symmetric CFTs with even power $m$ using the $\epsilon$ expansion, as one needs to be more careful about the quantization scheme and it is not an easy task to handle multidimensional Stokes sectors, i.e., Lefschetz thimbles. See Felski:2021evi and references therein. Furthermore, it seems more natural to use the combination $i\phi$ as the building block for $\mathcal{PT}$ symmetric theories. Recently, the low dimensional $\mathcal{PT}$ symmetric models have been revisited by bootstrap methods in Li:2022prn ; Khan:2022uyz ; Li:2023nip ; John:2023him ; Li:2023ewe ; Li:2024rod , using the equations of motion in the Lagrangian or Hamiltonian formalism. It would be interesting to extend the low dimensional bootstrap methods to higher dimensions. Along the lines of analytic continuation, one may also bootstrap the non-perturbative conformal field theory by the homotopy continuation method Li:2024ggr , where the $\epsilon$ expansion can provide regular boundary conditions near but away from the potentially dangerous Gaussian limit.  [^32] Conformal symmetry implies the existence of zero-norm states associated with specific Gaussian scaling dimensions. Usually, the zero-norm states should decouple to avoid inconsistency, such as divergences in the Gaussian limit.
In addition, the $\phi^{4}$ deformation of a canonical free theory with a boundary or monodromy/line defect has been studied using the multiplet recombination Yamaguchi:2016pbj ; Dey:2020jlc ; Nishioka:2022odm ; Nishioka:2022qmj . An interesting future direction is to consider the $\phi^{m}$ deformations and higher-derivative generalizations, as well as other types of defects. Another direction would be to study CFTs on nontrivial manifolds. The case of real projective space has been carried out for the $\phi^{3}$ and $\phi^{4}$ theory Hasegawa:2016piv ; Hasegawa:2018yqg . It would also be interesting to apply the multiplet recombination method to CFTs at finite temperature.
@@ Appendix C Representation theory of $S_{N+1}$
$\dim V_{(N-1,2)}$ $=\frac{(N+1)!}{N\times(N-1)\times(N-3)\times\ldots\times 1\times 2\times 1}=\frac{(N+1)(N-2)}{2}\,,$ (130)
@@
Comparing the dimensions with those of (128), (129), (130) and (131), we conclude that (132), (133), (134) and (135) correspond to the Young diagrams $(N+1)$, $(N,1)$, $(N-1,2)$ and $(N-1,1^{2})$. [^35] The dimension of (134) follows from the symmetric tracelessness and the constraints $\sum_{a,b=1}^{N}T_{iab}\sum_{c,d=1}^{N}t^{(3)}_{abcd}\phi_{c}\phi_{d}=0\,,$ (137) where $i=1,2,\ldots,N$. From these examples, we see that the tensor product decomposition resembles that in the O($N$) model, except that here the symmetric traceless part further decomposes into two irreducible representations $V_{(N,1)}$ and $V_{(N-1,2)}$. Since $V_{(N,1)}$ is the standard representation, we use the term “symmetric traceless representation” to refer to $V_{(N-1,2)}$. Moreover, we refer to $V_{(N+1)}$ and $V_{(N-1,1^{2})}$ as the “trivial representation” and the “antisymmetric representation”. [^36] The antisymmetric representation $V_{(N-1,1^{2})}$ is the exterior power $\Lambda^{2}V_{(N,1)}$ of the standard representation $V_{(N,1)}$.
@@ Appendix D Tensor structures
The tensor product decomposition of $(N,1)^{\otimes 4}$ contains the trivial representation four times, so there are four linearly independent rank-4 invariant tensors wallace1978spin . Using $e_{a}^{\alpha}$ as defined in (126), we identify four rank-4 $S_{N+1}$-invariant tensors: $\delta_{ab}\delta_{cd}$, $\delta_{ac}\delta_{bd}$, $\delta_{ad}\delta_{bc}$ and $\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}$. We then use these tensors to define a basis for the four-point function $\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\phi_{c}(x_{3})\phi_{d}(x_{4})\rangle$:
@@ Appendix F An OPE coefficient in generalized $\phi^{2n+1}$ theory ($n>1$)
Appendix G The $N\rightarrow 0$ limit, logarithmic CFTs and logarithmic recombination
As mentioned in footnote 8, the $N\rightarrow 0$ limit of the Potts model corresponds to a logarithmic CFT. A key feature is that the tensor structure (140) diverges in the $N\rightarrow 0$ limit. This singularity can be resolved by mixing bilinear operators. We will see below that this mixing leads to logarithmic structures and nonvanishing off-diagonal two-point functions. This means that the primaries are not independent and form logarithmic multiplets, which can be viewed as a logarithmic version of the multiplet recombination. The discussion in this appendix applies to the canonical case $k=1$ and the higher-derivative cases with $k>1$.
In the above, the tensor structure (140) is associated with the two-point functions of bilinear operators $\mathcal{J_{\text{ST}}}$ in the symmetric traceless (ST) representation. To cancel the $1/N$ pole, $\mathcal{J_{\text{ST}}}$ should mix with $\mathcal{J_{\text{trivial}}}$ and they form new operators.373737 Although the tensor structure (140) is also singular in the $N\rightarrow-1$ limit, the two-point function of $\phi^{2}_{\text{ST}}$ vanishes due to the normalization (see (162)). At order $\epsilon^{0}$, the OPE coefficient vanishes as well (see (169)), but it can be made finite by a singular change in the normalization $\bar{\phi}^{2}=\frac{1}{N+1}\phi^{2}_{\text{ST}}$. For instance, the two-point functions of the spin-0 operators with $m=0$ take the form
$\langle\phi^{2}_{\text{trivial}}(x_{1})\phi^{2}_{\text{trivial}}(x_{2})\rangle$ $=\frac{2N}{x_{12}^{\Delta_{\smash{\phi^{2}_{\text{trivial}}}}}}\,,$ (161)
$\langle\phi^{2}_{ab,\text{ST}}(x_{1})\phi^{2}_{cd,\text{ST}}(x_{2})\rangle$ $=\frac{2(N+1)^{2}(N-1)^{2}}{x_{12}^{\Delta_{\smash{\phi^{2}_{\text{ST}}}}}}t^{(3)}_{abcd}\,.$ (162)
Here the normalization of the two-point functions is the same as that in the free theory, i.e., the two-point functions are given by Wick’s theorem and $\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\rangle=|x_{1}-x_{2}|^{-2\smash{\Delta_{\phi}}}\delta_{ab}$. The explicit expressions for $\phi^{2}_{ab,\text{ST}}$ and $\phi^{2}_{\text{trivial}}$ can be found in (132) and (134). A linear combination that is regular in the $N\rightarrow 0$ limit reads
$\bar{\phi}^{2}_{ab}=\phi^{2}_{ab,\text{ST}}-\frac{\delta_{ab}}{N}\phi^{2}_{\text{trivial}}\,.$ (163)
The coefficient $-1/N$ ensures that the correlation functions $\langle\bar{\phi}^{2}_{ab}(x_{1})\bar{\phi}^{2}_{cd}(x_{2})\rangle$ and $\langle\phi_{a}(x_{1})$ $\phi_{b}(x_{2})\bar{\phi}^{2}_{cd}(x_{3})\rangle$ are regular in the $N\rightarrow 0$ limit. This coefficient is not unique because one can add more $N$-dependent terms as long as they are not singular in the $N\rightarrow 0$ limit. In our convention (163), the two-point function of $\bar{\phi}^{2}_{ab}$ reads
$\lim_{N\rightarrow 0}\langle\bar{\phi}^{2}_{ab}(x_{1})\bar{\phi}^{2}_{cd}(x_{2})\rangle$
=\;= $\lim_{N\rightarrow 0}\Bigg{[}\frac{1}{x_{12}^{2\Delta_{\smash{\phi^{2}_{\text{ST}}}}}}\Bigg{(}2\left(N^{2}+N-1\right)\delta_{ab}\delta_{cd}+(N+1)^{2}(N-1)^{2}\left(\delta_{ac}\delta_{bd}+\delta_{ad}\delta_{bc}\right)$
$\hskip 30.00005pt-2(N+1)(N-1)\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}+\frac{8\epsilon}{7-3N}\log x_{12}\,\delta_{ab}\delta_{cd}+O(\epsilon^{3/2})\Bigg{)}\Bigg{]}$
=\;= $\frac{1}{x_{12}^{2\Delta_{0}}}\left(-2\delta_{ab}\delta_{cd}+2\left(\delta_{ac}\delta_{bd}+\delta_{ad}\delta_{bc}\right)+2\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}+\frac{8\epsilon}{7}\log x_{12}\,\delta_{ab}\delta_{cd}+O(\epsilon^{3/2})\right)\,.$ (164)
In the last line, $\Delta_{0}$ denotes the degenerate scaling dimension
$\Delta_{0}=\Delta_{\smash{\phi^{2}_{\text{ST}}}}\big{|}_{N\rightarrow 0}=\Delta_{\smash{\phi^{2}_{\text{trivial}}}}\big{|}_{N\rightarrow 0}\,,$ (165)
which can be checked to order $\epsilon^{1}$ using (69). The logarithm arises due to the $\epsilon$ expansion of $x_{12}^{2\Delta_{\smash{\phi^{2}_{\text{ST}}}}-2\Delta_{\smash{\phi^{2}_{\text{trivial}}}}}$, which comes from the contribution of $\langle\phi^{2}_{\text{trivial}}(x_{1})\phi^{2}_{\text{trivial}}(x_{2})\rangle$. Here we have used (69).
Apart from the regularity of the two-point function $\langle\bar{\phi}^{2}_{ab}(x_{1})\bar{\phi}^{2}_{cd}(x_{2})\rangle$, the combination (163) also lead to nonvanishing off-diagonal two-point functions:
$\lim_{N\rightarrow 0}\begin{pmatrix}\langle\bar{\phi}^{2}_{ab}(x_{1})\bar{\phi}^{2}_{cd}(x_{2})\rangle&\delta_{cd}\langle\bar{\phi}^{2}_{ab}(x_{1})\phi^{2}_{\text{trivial}}(x_{2})\rangle\\ \delta_{ab}\langle\phi^{2}_{\text{trivial}}(x_{1})\bar{\phi}^{2}_{cd}(x_{2})\rangle&\;\delta_{ab}\delta_{cd}\langle\phi^{2}_{\text{trivial}}(x_{1})\phi^{2}_{\text{trivial}}(x_{2})\rangle\end{pmatrix}$
=\;= $\frac{1}{x_{12}^{2\Delta_{0}}}\begin{pmatrix}\bar{t}_{abcd}+\frac{8\epsilon}{7}\log x_{12}\,t^{(1)}_{abcd}&\;-2t^{(1)}_{abcd}\\ -2t^{(1)}_{abcd}&0\end{pmatrix}+O(\epsilon^{3/2})\,,$ (166)
where $t^{(1)}_{abcd}$ is given by (138) and $\bar{t}_{abcd}$ is defined as
$\bar{t}_{abcd}=-2\delta_{ab}\delta_{cd}+2\left(\delta_{ac}\delta_{bd}+\delta_{ad}\delta_{bc}\right)+2\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}\,.$ (167)
We interpret the matrix elements as the two-point functions of the primaries in a rank-2 logarithmic multiplet. Although the norm of $\phi^{2}_{\text{trivial}}$ vanishes in the $N\rightarrow 0$ limit, $\phi^{2}_{\text{trivial}}$ is not a null state and have nonzero correlation functions with other operators.
Let us verify that the three-point function $\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\bar{\phi}^{2}_{cd}(x_{3})\rangle$ is regular as well. We consider the three-point functions of two $\phi$’s and a bilinear operator
$\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\phi^{2}_{\text{trivial}}(x_{3})\rangle$ $=\left(2+O(\epsilon^{1/2})\right)\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}\delta_{ab}\,,$ (168)
$\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\phi^{2}_{cd,\text{ST}}(x_{3})\rangle$ $=\left(2(N+1)(N-1)+O(\epsilon^{1/2})\right)\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}t^{(3)}_{abcd}\,.$ (169)
Here $h_{ijk}=\Delta_{i}+\Delta_{j}-\Delta_{k}$, and $\Delta_{i}$ is the scaling dimension of the operator at position $x_{i}$. The leading term in the OPE coefficient for $\phi^{2}_{\text{trivial}}$ does not depend on $N$, so it remains finite in the $N\rightarrow 0$ limit, as mentioned above. However, the tensor structure $t^{(3)}_{abcd}$ in the second line diverges as $N$ approaches zero. This singular behaviour is also resolved by the linear combination (163), yielding the three-point function
$\lim_{N\rightarrow 0}\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\bar{\phi}^{2}_{cd}(x_{3})\rangle=\;$ $\left(-\bar{t}_{abcd}+O(\epsilon^{1/2})\right)\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}\Bigg{|}_{N\rightarrow 0}\,.$ (170)
Here $\Delta_{3}$ is understood as $\Delta_{\smash{\phi^{2}_{\text{ST}}}}$.
According to the results for the correlation functions, the two independent primaries $\phi^{2}_{\text{trivial}}$ and $\phi^{2}_{\text{ST}}$ mix and form the logarithmic primary $\bar{\phi}^{2}_{\text{ab}}$ in the limit $N\rightarrow 0$. The two primary states associated with $\bar{\phi}^{2}_{\text{ab}}$ and $\phi^{2}_{\text{trivial}}$ are not orthogonal to each other, giving rise to nonvanishing off-diagonal two-point functions. In other words, $\{\phi^{2}_{\text{trivial}}\}$ and $\{\bar{\phi}^{2}_{\text{ab}}\}$ recombine into a logarithmic multiplet. As the two multiplets are related by the derivative with respective to the scaling dimension $\Delta$, the logarithmic recombination is different from the multiplet recombination based on the derivative with respective to $x^{\mu}$ in the main text. The logarithmic version of the multiplet recombination may deserve further investigations.
Furthermore, the degeneracy (165) extends to all bilinear operators in the trivial and symmetric traceless representations. Using (69), we obtain
$\Delta_{\mathcal{J}_{\text{trivial}}}\big{|}_{N\rightarrow 0}=\Delta_{\mathcal{J}_{\text{ST}}}\big{|}_{N\rightarrow 0}=d-2k+2m+\ell$
$+\left(\frac{1}{3+\frac{(-1)^{k}k!(3k)_{k}}{2^{4k-5}(1/2)_{k}^{2}}}+\frac{(2k)!(2k-m)_{k+m}}{\left(8k!+\frac{3\times 2^{4k-2}(-1)^{k}(1/2)_{k}^{2}}{(3k)_{k}}\right)(1-3k-\ell-m)_{2k}(-m)_{m}}\right)\epsilon+O(\epsilon^{3/2})\,.$ (171)
These equations hold for arbitrary spin $\ell$, so the entire Regge trajectories associated with $\mathcal{J}_{\text{trivial}}$ and $\mathcal{J}_{\text{ST}}$ with the same $m$ are degenerate. The spinning operators should also exhibit logarithmic structures. To order $\epsilon^{1}$, this agrees with the fact that the $k=1$ crossover exponent is one wallace1978spin :
$\frac{d-\Delta_{\phi^{2}_{\text{ST}}}|_{N\rightarrow 0}}{d-\Delta_{\phi^{2}_{\text{trivial}}}|_{N\rightarrow 0}}=1+O(\epsilon^{3/2})\,.$ (172)
Note that (G) generalizes this relation to $k>1$ and arbitrary spin, to leading order in $\epsilon$.
There is a similar story in the $N\rightarrow 0$ limit of the generalized O($N$)-symmetric $\phi^{4}$ theory. For general $N$, the scalar two-point functions of singlet (S) and symmetric traceless (ST) bilinear operators are
$\langle\phi^{2}_{\text{S}}(x_{1})\phi^{2}_{\text{S}}(x_{2})\rangle$ $=\frac{2N}{x_{12}^{\Delta_{\phi_{\text{S}}}}}\,,$ (173)
$\langle\phi^{2}_{ab,\text{ST}}(x_{1})\phi^{2}_{cd,\text{ST}}(x_{2})\rangle$ $=\frac{2}{x_{12}^{\Delta_{\phi_{\text{ST}}}}}\left(\frac{\delta_{ac}\delta_{bd}+\delta_{ad}\delta_{bc}}{2}-\frac{1}{N}\delta_{ab}\delta_{cd}\right)\,,$ (174)
where $\phi^{2}_{\text{S}}=\sum_{i=1}^{N}\phi_{i}\phi_{i}$ and $\phi^{2}_{ab,\text{ST}}=\phi_{a}\phi_{b}-\frac{\delta_{ab}}{N}\sum_{i=1}^{N}\phi_{i}\phi_{i}$. To avoid the singularity in the $N\rightarrow 0$ limit, the operator $\phi^{2}_{\text{ST}}$ should mix with $\phi^{2}_{\text{S}}$ (see e.g. Hogervorst:2016itc ). A regular combination is
$\bar{\phi}^{2}_{ab}=\phi^{2}_{ab,\text{ST}}+\frac{\delta_{ab}}{N}\phi^{2}_{\text{S}}\,.$ (175)
As in (163), the coefficient in front of $\phi^{2}_{\text{S}}$ is not unique. Furthermore, we expect that the degeneracies extend to the spinning operators. The scaling dimensions of the singlet and symmetric traceless bilinear operators indeed become identical in the $N\rightarrow 0$ limit:
$\Delta_{\mathcal{J}_{\text{S}}}\big{|}_{N\rightarrow 0}=\Delta_{\mathcal{J}_{\text{ST}}}\big{|}_{N\rightarrow 0}$
=\;= $d-2k+2m+\ell-\frac{(k-1)!(k+1)_{k}^{2}}{128(k)_{k-m}}\left(\frac{1}{(-k)_{k}(2k-m)_{k+m}}+\frac{3}{(-m)_{m}(m+\ell)_{2k}}\right)\epsilon^{2}+O(\epsilon^{3})\,.$ (176)
As in (G), the spin $\ell$ is arbitrary, implying that the entire Regge trajectories are degenerate in the limit $N\rightarrow 0$ for each $m$, giving rise to logarithmic structures as in the Potts model.